OpenStudy (anonymous):

Write the rule for the following arithmetic sequence: 17, 22, 27, 32, …

OpenStudy (anonymous):

a,b,c,d,e,f,... b-a=5 c-b=5 d-c=5 ... 22-17=5 27-22=5 32-27=5 after 32 is 37 because 37-32=5 do you realize?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

listen to me...22-17=? say...

OpenStudy (anonymous):

@ymferreira